Simplify.
-7[27-(12-39)]=
Is the answer to this 273-189-84 ??
I a close or way off the planet??

Answer by rapaljer(4671)

I don't know where you got those numbers. It should be like this. Do what is inside first, like 12-39 = -27
-7[27-(12-39)]=
-7[27-(-27)]=
-7[27+27]
-7*54
-378

R^2